# DAIRstega: Dynamically Allocated Interval-Based Generative Linguistic Steganography with Roulette Wheel

### v4

@@ Abstract
Linguistic steganography (LS) tasks aim to generate steganographic text (stego) based on secret. Only authorized receivers can perceive and extract secrets, thereby protecting privacy. However, existing generative LS schemes often do not consider the conditional probability of tokens in the candidate pool, and allocate one or the same number of codings to all tokens. The tokens with lower probabilities are selected to embed secrets that will affect the quality of stegos. As a result, the stegos are easy to perceive and detect. This paper proposes the LS scheme based on dynamically allocated intervals, called DAIRstega. DAIRstega uses the idea of the roulette wheel and takes the conditional probabilities of tokens as the main basis for allocating the roulette area (i.e., the interval length). Thus, the token with a larger conditional probability is allocated more. The secret will be more likely to select the tokens with larger probabilities. In the allocation process, we design some functions between probability and allocated interval length. Based on the invisible characteristics of LS, we give three constraints that need to be met to design the function. To simplify the form, the expression of the allocation function is condensed. Furthermore, DAIRstega can receive additional instruction and controllably generate stegos. Rich experiments show that the proposed embedding way and DAIRstega perform superior to the existing ways and LS schemes, which shows strong perceptual, statistical, and semantic concealment and anti-steganalysis ability. This scheme can also generate high-quality longer stegos, improving the deficiencies in this task. The experiment also verified that DAIRstega can be used as a secure watermarking scheme, providing some ideas for its development.
@@ 1 Introduction
Steganography, a technology of information hiding (Shannon 1949), embeds secrets within media such as texts (Yang et al. 2023, 2021) and images, obtaining steganographic media that is sensory indistinguishable from normal media. Only authorized receivers can perceive whether the media is steganographic and accurately extracts the secret. Benefiting from the ability of text to be transmitted losslessly over public channels, research on Linguistic Steganography (LS) has explosive growth (Lu et al. 2023; Wang et al. 2023a; Yang et al. 2023). To increase the anti-detection ability of steganographic text (stego), the focus of work has shifted from modified LS (Huo and Xiao 2016) represented by synonym replacement (Kim, Zaiane, and Goebel 2010) to generative LS based on secret (Li et al. 2021; Yang et al. 2021).
@@
The concealment of stego determines the success of covert communication to a great extent. Depending on the constraints, the concealment is manifested in three aspects: perceptual, statistical, and semantic. “Perceptual concealment” focuses on ensuring stegos with complement and fluency. “Statistical concealment” requires the distribution of stegos to be close to that of the natural texts (covers). “Semantic concealment” aims to stego that is coherent and semantically controllable. The concealment of stego mainly depends on the design of the steganography embedding way. An excellent embedding way ensures that the conditional probability (CP) distribution of the text is not changed as much as possible while embedding the secret. In this way, stegos are difficult to perceive and detect by unauthorized receivers. Most existing LS schemes have made efforts in this regard. Yang et al. (Yang et al. 2019) and Ding et al. (Ding et al. 2023a) encoded each token in the candidate pool with Fixed length coding (FLC) and only use secret information to determine the next token to achieve the purpose of embedding. Yang et al. (Yang et al. 2021) and Zhou et al. (Zhou et al. 2021) designed the Huffman coding-based (HC) embedding, which generates stegos with good completeness and fluentness, improving perceptual concealment. Zhang et al. (Zhang et al. 2021) and Wang et al. (Wang et al. 2023a) proposed the Adaptive dynamic grouping coding-based (ADG) embedding. They achieved strong statistical concealment. Lu et al. (Lu et al. 2023) used contextual learning and variable length coding-based (VLC) to generate stegos. It improves semantic concealment.
@@
However, the embedding ways of almost all existing schemes allocate one or the same number of codings to the tokens in the candidate pool. That is, when embedding, the probability of selecting all tokens is the same. They do not consider that the CP of a portion of the tokens is not high. If the secret is determined to these tokens, the quality of stego generation will be dropped. This will affect the quality and statistical distribution, and thus the anti-detection ability. The results of the existing embedding ways are given in Figure 1. Therefore, it is necessary to explore a difficult-to-perceive and detect scheme with concealed embedding.
@@
Figure 1: The result of large language model (LLM) + embedding ways. $\uparrow$” and $\downarrow$” represent the higher / lower the value, the better the result. The metrics can be found in “Section 3.1”. The values can be seen in Table 1.
@@
This paper proposes the dynamically allocated interval-based LS scheme, called DAIRstega. Different from the existing LS, DAIRstega introduces the idea of the roulette wheel to design the secret embedding way. We believe that the quality of the tokens in the candidate pool is different. Thus, the tokens need to be given different shoot probabilities by secret, making it more likely that the secret will select tokens with higher CPs. Then, we explore the characteristics of the CP of tokens in the candidate pool, and design a variety of relationships between CPs and the number of allocated codings. By analyzing the adaptability of these allocated functions to the invisible information-hiding tasks of LS, the allocation expression is condensed. In addition, DAIRstega can receive instruction and more directly control the generation of the stego contents. Experiments show that DAIRstega has strong concealment and controllability.
The main contributions are summarized as follows.
@@
  • To augment the concealment of stegos, DAIRstega introduces the idea of the roulette wheel, and dynamically allocates the number of codings. The multiple functions between CPs and allocated numbers are also given. According to Occam’s razor, the functions are condensed.
  • To ensure the adaptability of the functions designed, we gave three constraints that need to be met.
@@
  • To verify the effectiveness of the proposed embedding way and DAIRstega, we performed experiments regarding various concealment and anti-steganalysis.
  • Experiment verified that DAIRstega can also be regarded as a secure watermarking scheme with high fidelity, consistency, and abuse-resistant.
@@ 2.1 Modeling
The covert communication hopes to avoid Eve’s perception and detection and needs to satisfy the following:
@@
where, ${P_{C}}$ and ${P_{S}}$ are the distribution of cover and stego, $d(\cdot)$ is the difference, $\varepsilon$ is a very small value greater than 0. $\sim$ is following a certain distribution. The above assumptions cover most situations where the secret is transmitted. Complying with the Formula 1 can improve the concealment and ensure the conduct of covert communication. The results of this aspect are shown in Table 4 and Table 5.
@@ 2.2 DAIRstega Overall
This scheme can generate stegos that are strongly concealed and controllable. It can satisfy the extension of covert communication, i.e. Formula 1. To make the generated stegos more consistent with the CP distribution of covers, DAIRstega uses the idea of the roulette wheel to dynamically allocate tokens in the candidate pool to coding intervals of different lengths. That is, the larger the CP is, the larger the roulette area will be allocated, and the greater the shoot that this token will be selected by secret. To make the generated stegos controllable, we introduce the Prompt, taking instructions to control generate. Instruction can be public topics, posted questions, or other requirements in social networks. Figure 3 shows the DAIRstega’s framework.
Figure 3: The DAIRstega’s framework. The “stego generation process” consists of two parts: the “CP generation” and “Embedding” modules. Different from existing LS works, DAIRstega uses the idea of the non-uniform roulette wheel to dynamically allocate different roulette areas (numbers of codings) to tokens in the candidate pool. In this way, the secret is more likely to fall on the tokens with larger CPs, ensuring the quality of stegos. In addition, DAIRstega can not only receive secret information (in the form of bitstreams), but also input the instruction, and finally generate stego that conforms to the instruction.
@@ 2.3 Details
Stego generation process.
It is composed of the “CP generation” ($PG$) and “Embedding” ($Em$) modules. The process formulas are as follows:
$\left\{\begin{array}[]{l}PG:LM(I)\Rightarrow\mathbf{P}\\ Em:\text{locate}\;\{{m_{t}}\;\text{in}\;DAI(BC(\mathbf{P}))\}\Rightarrow i,\\ \quad\quad\;\ \text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ i=Em(PG(I))\end{array}\right.,$ (2)
@@
Embedding module.
To ensure that the embedding does not excessively affect the quality of the generated stegos. In our work, low-CP tokens should not be added to the candidate pool. We sort and intercept the first top_k tokens with higher CPs to build the candidate pool. The $\mathbf{P}$ normalization $\mathbf{P}^{\prime}=[{p^{\prime}_{1}},\cdots,{p^{\prime}_{top\_k}}]$ of tokens in candidate pool is:
@@
Further exploration and discovery, if the linear function is used, the embedding rate can be improved. Tokens with very high $p$ will be allocated to most encodings, and a large interval length will result in a short prefix of this interval regardless of how many bits are read at a time. Longer stego needs to be generated to embed the given bits, resulting in a lower embedding rate (unit: bit per word). So we try to design the nonlinear functions to improve the embedding rate.
Using logarithmic and exponential operations, a variety of nonlinear allocations can be designed. However, not all nonlinear allocations are suitable for LS information hiding. We define three constraints that the functions need to be met:
Constraint 1:
${\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}\geq\frac{c}{1-x},c\in\mathbb{R}^{+},\forall x\in\left[\epsilon^{+},1\right),$ (5)
@@
where, $x$ is a variable that meets all CPs. $\epsilon^{+}$ is a smaller value greater than 0, such as 0.1, 0.2, and so on. Although the CP distribution is discrete, the independent variable of the distribution function is continuous, so it is expressed here using derivatives. This constraint shows that the number of allocations needs to increase with the increase of CP, and ensures that the change rate of the allocated number is not too small. If ${\textstyle{{\partial\;DAI(x)}\over{\partial\;x}}}\leq\varepsilon$, which means the number of allocation tokens is similar. It will degenerate into a similar allocation of the same number of codings as the existing work, thus increasing the selection of a poor token.
Constraint 2:
${\frac{{\partial^{2}{DAI}{(x)}}}{{\partial x^{2}}}}\leq 0,\forall x\in\left[\epsilon^{+},1\right).$ (6)
When the condition is equal, the corresponding function will be transformed into a linear function. This constraint indicates that the entire allocation function should be an increasing function with decreasing growth. If the function has an increasing growth trend, then tokens with larger CPs will be allocated to larger intervals, and the same prefixes at the beginning and end of the interval will be shortened. Thus, the embedding rate will be decreased.
Constraint 3:
$DAI(x)\geq x\;\wedge\;{\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}-{\frac{{\partial\ x}}{{\partial\ x}}}\leq 0,\forall x\in\left[\epsilon^{+},1\right).$ (7)
This constraint shows that the number of codings allocated to tokens with smaller CP should be appropriately increased to increase their probability of being selected. This can also suppress the problem of too large intervals of larger CP to a certain extent.
Here, we give several nonlinear allocation functions:
@@
The smaller $\beta$ is, the smaller the interval of tokens with larger CP will be, while still satisfying the premise that the interval of tokens with larger CP is larger. It will increase the length of the prefix, while reducing the length of stegos to be generated, improving the embedding rate.
Please note that these functions are not fixed, here are just a few examples that meet the constraints. The constraints we proposed are not the most perfect, and the final condensed expression is also for simplicity and easy to understand, and is not the best. This can just provide thinking for subsequent research that is more conducive to hiding information.
Loop the stego generation process, the stego will be obtained. Algorithm 1 illustrates the embedding process.
@@
Secret extraction process.
Bitstream extraction and embedding are a pair of inverse operations. Same as Discop (Ding et al. 2023b), we also need to use the same scheme, initial context (instruction), and stegos. The instructions remain requirements visible to everyone in social networks. The formula of the extraction process $Ex$ is as follows:
$\left\{\begin{array}[]{l}Ex:\text{locate}\;\ \{{i_{n+1}}\;\ {\rm{in}}\;\ DAI(BC(P))\},\\ \quad\quad\;\text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ B_{j}=Ex(PG(I)),B_{j}\in B\end{array}\right..$ (10)
The interval is located by the stego’s next token at this moment, and the same prefix is the current extracted secret.

### v5

# Dynamically Allocated Interval-Based Generative Linguistic Steganography with Roulette Wheel

@@ Abstract
Existing linguistic steganography schemes often overlook the conditional probability (CP) of tokens in the candidate pool, allocating the one coding to all tokens, which results in identical selection likelihoods. This approach leads to the selection of low-CP tokens, degrading the quality of stegos and making them more detectable. This paper proposes a scheme based on the interval allocated, called DAIRstega. DAIRstega first uses a portion of the read secret to build the roulette area. Then, this scheme uses the idea of the roulette wheel and takes the CPs of tokens as the main basis for allocating the roulette area (i.e., the interval length). Thus, tokens with larger CPs are allocated more area. The secret will have an increased likelihood of selecting a token with a higher CP. During allocation, we designed some allocation functions and three constraints to optimize the process. Additionally, DAIRstega supports prompt-based controllable generation of stegos. Rich experiments show that the proposed embedding way and DAIRstega perform better than the existing ways and baselines, which shows strong perceptual, statistical, and semantic concealment, as well as anti-steganalysis ability. It can also generate high-quality longer stegos, addressing the deficiencies in this task. DAIRstega is confirmed to have potential as a secure watermarking, offering insights for its development. Our codes and data are available at: https://github.com/WangYH-BUPT/DAIRstega.
keywords:
Linguistic steganography , Embedding way , Allocation funtion , Interval allocated , Roulette wheel , Large language models , Concealment
t1t1footnotetext: The paper has been accepted by Applied Soft Computing.t1t1footnotetext: Thanks for the support provided by MindSpore Community.t2t2footnotetext: This work is supported by the National Natural Science Foundation of China (Grant U21B2020) and supported by BUPT Excellent Ph.D. Students Foundation (Grant CX2023120, CX20241055).
@@ 1 Introduction
Steganography, a technology of information hiding [1], embeds secrets within media such as texts [2, 3] and images, obtaining steganographic media that is perceptually indistinguishable from normal media. Only authorized receivers can perceive whether the media is steganographic and accurately extract the secrets, thereby protecting privacy. Benefiting from the ability of text to be transmitted losslessly over public channels, research on linguistic steganography has explosive growth [4, 5, 2]. To increase the anti-detection of steganographic text (stego), the focus of work has shifted from modified schemes [6] represented by synonym replacement [7] to generative schemes [8, 3].
@@
The concealment of stego determines the success of covert communication to a great extent. Depending on the constraints, the concealment is manifested in three aspects: perceptual, statistical, and semantic. “Perceptual concealment” focuses on ensuring stegos with complement and fluency. “Statistical concealment” requires the distribution of stegos to be close to that of the natural texts (covers). “Semantic concealment” aims to stego that is coherent and semantically controllable. The concealment of stego mainly depends on the design of the steganography embedding way. An excellent embedding way ensures that the conditional probability (CP) distribution of the text is not changed as much as possible while embedding the secret. In this way, stegos are difficult to perceive and detect by unauthorized receivers. Most existing schemes have made efforts in this regard. Yang et al. [9] and Ding et al. [10] encoded each token in the candidate pool with Fixed length coding (FLC) and only used secret information to determine the next token to achieve the purpose of embedding. Yang et al. [3] and Zhou et al. [11] designed the Huffman coding-based (HC) embedding, which generates stegos with good completeness and fluency, improving perceptual concealment. Zhang et al. [12] proposed the Adaptive dynamic grouping coding-based (ADG) embedding. They achieved strong statistical concealment. Kaptchuk et al. [13] and Wang et al. [5] proposed the Arithmetic coding-based (AC) embedding. Lu et al. [4] used contextual learning and variable length coding-based (VLC) to generate stegos. It improves semantic concealment.
@@
However, the embedding ways of almost all existing schemes allocate one or the same number of codings to the tokens in the candidate pool. That is, when embedding, the likelihood of selecting all tokens is the same. They do not consider that the CP of a portion of the tokens is not high. If the secret is determined to these tokens, the quality of stego generation will be dropped. The example is shown in the Figure 1.
Figure 1: Examples of existing embedding ways and the proposed embedding way. After selecting $n$ tokens with larger CPs to build a candidate pool, the existing way assigns a coding to each token in the candidate pool. Since the secret binary can be arbitrary, to embed the secret, the existing embedding way selects all tokens with the same likelihood of $1/n$. The proposed way will build a roulette area and allocate it according to the CP value to ensure that the larger the CP of the token, the larger the allocation area. Therefore, the proposed way has a higher likelihood of selecting a better token. In the end, the quality of the token determined by the proposed way is generally better than that of the existing ways.
Tokens with low CPs will reduce the quality of stego and cause significant differences in its statistical characteristics from cover. Since steganalysis detects stegos by capturing these differences [14, 15], poor-quality stego is easier to identify and has weaker anti-steganalysis capabilities. The performance of stegos obtained by different embedding ways is shown in Figure 2. Therefore, it is necessary to explore a difficult-to-perceive and detect scheme with concealed embedding.
@@
Figure 2: The result of large language model (LLM) + embedding ways. $\uparrow$” and $\downarrow$” represent the higher / lower the value, the better the result. The metrics are found in “Section 4.1”. The values are seen in Table 2.
@@
This paper proposes the dynamically allocated interval-based steganography, called DAIRstega. Different from the existing schemes, DAIRstega introduces the idea of the roulette wheel to design the embedding way. We believe that the quality of the tokens in the candidate pool is different. Thus, the tokens need to be given different shoot likelihoods by secret, increasing the likelihood of selecting tokens with higher CP. Then, we explore the characteristics of the CP of tokens, and design a variety of relationships between CPs and the number of allocated codings. By analyzing the adaptability of these allocated functions to the invisible information-hiding tasks, the allocation expression is condensed. In addition, DAIRstega can receive instructions and directly control the stego generation.
DAIRstega is built upon the MindSpore framework and the contributions are summarized as follows.
  • 1.
@@
    To augment the concealment of stegos, DAIRstega introduces the idea of the roulette wheel, and dynamically allocates the number of codings. The functions between CPs and allocated numbers are also given. According to Occam’s razor, the functions are condensed.
  • 2.
    To ensure the adaptability of the functions designed, we gave three constraints.
  • 3.
@@
  • 4.
    To verify the effectiveness of the proposed embedding way and DAIRstega, we performed experiments regarding perceptual, statistical, and semantic concealment, as well as anti-steganalysis performance.
  • 5.
    Experiment verified that DAIRstega can also be regarded as a secure watermarking scheme with high quality, consistency, and unforgeability.
2 Related work
Perceptual concealment” focuses on ensuring that the scheme generates stego with complete and fluent. This is the most basic requirement of steganography. In pursuit of this objective, Fang et al. [16] proposed an LSTM-based steganography scheme. This scheme segments the vocabulary into several sets based on bit blocks. It then selects tokens with the highest probability that corresponds to the secret information from the candidate pool. Ding et al. [10] combined the conditional generation strategy with the replacement technique, using text sequences as auxiliary data in the stego-generation process to enhance the embedding capabilities. Yang et al. [9] trained a language model to generate stegos. Using HC-based (Huffman coding) and FLC-based (Fixed length coding) to encode the candidate pool, these stegos have excellent completeness and fluentness.
Statistical concealment” requires the stegos distribution to be close to the covers. To achieve this goal, Yang et al. [3] designed an encoder-decoder structure, and VLC-based (Variable length coding) is used to encode the CP. To further reduce the distribution difference between cover and stego, Zhang et al. [12] used adaptive dynamic grouping coding to embed the secret recursively. Zhou et al. [11] used a generative adversarial network (GAN) to design an adaptive probability distribution steganography. These ensure statistical concealment.
Semantic concealment” aims to generate a stego that is coherent and semantically controllable. To this end, Li et al. [8] put forward a steganography scheme based on the knowledge graph. Yang et al. [2] utilized semantic information encoding to embed secret information, realizing the effect of maintaining semantics and increasing the embedding capacity during the translation process. Wang et al. [5] leveraged the relevance of social network context to enhance contextual semantic relevance while maintaining embedding rates. Lu et al. [4] used contextual learning and GPT2 guidance to generate stegos.
@@ 3.1 Modeling
The covert communication hopes to avoid Eve’s perception and detection and needs to satisfy:
@@
where, ${P_{C}}$ and ${P_{S}}$ are the distribution of cover and stego, $d(\cdot)$ is the difference, $\varepsilon$ is a very small value greater than 0. $\sim$ is following a certain distribution. The above assumptions cover most situations where the secret is transmitted. Complying with the Formula 1 can improve the concealment and ensure the conduct of covert communication. The results of this aspect are shown in Table 5.
@@ 3.2 DAIRstega Overall
This scheme can generate stegos that are strongly concealed and controllable. It can satisfy the extension of covert communication, i.e., Formula 1. To make the generated stegos more consistent with the CP distribution of covers, DAIRstega uses the idea of the roulette wheel to allocate tokens dynamically in the candidate pool to coding intervals of different lengths. That is, the larger the CP is, the larger the roulette area will be allocated, and the greater the shoot that this token will be selected by secret. Instructions can be public topics, posted questions, or other requirements. Figure 4 shows the DAIRstega’s framework.
Figure 4: The DAIRstega’s framework. The “stego generation process” consists of two parts: the “CP (conditional probability) generation” and “Embedding” modules. Different from existing works, DAIRstega uses the idea of the non-uniform roulette wheel to dynamically allocate different roulette areas (numbers of codings) to tokens in the candidate pool. In this way, the secret is more likely to fall on the tokens with larger CPs, ensuring the quality of stegos. In addition, DAIRstega can not only receive secret information (in the form of bitstreams), but also input the instruction, and finally generate stego that conforms to the instruction.
@@ 3.3 Details
Stego generation process
It is composed of the “CP (conditional probability) generation” ($CPG$) and “Embedding” ($Em$) modules. The process formulas are as follows:
$\left\{\begin{array}[]{l}CPG:LM(I)\Rightarrow\mathbf{P}\\ Em:\text{locate}\;\{{m_{t}}\;\text{in}\;DAI(BC(\mathbf{P}))\}\Rightarrow i,\\ \quad\quad\;\ \text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ i=Em(CPG(I))\end{array}\right.,$ (2)
@@
Embedding module
To ensure that the embedding does not excessively affect the stego quality. In our work, low-CP tokens should not be added to the candidate pool. We sort and intercept the first top_k tokens with higher CPs to build the candidate pool. The $\mathbf{P}$ normalization $\mathbf{P}^{\prime}=[{p^{\prime}_{1}},\cdots,{p^{\prime}_{top\_k}}]$ of tokens in candidate pool is:
@@
Further exploration and discovery, if the linear function is used, the payload (unit: bit per word) can be improved. Tokens with very high $p$ will be allocated to most codings, and a large interval length will result in a short prefix of this interval regardless of how many bits are read at a time. Longer stego needs to be generated to embed the given bits, resulting in a lower payload. So, we try to design the nonlinear functions to improve the payload. Using logarithmic and exponential operations, a variety of nonlinear allocations can be designed. However, not all nonlinear allocations are suitable for information hiding. We define three constraints that the functions need to meet. $x$ is a variable that meets all CPs. Since the CP value of the token in the candidate pool is relatively large, $\epsilon^{+}$ is a smaller value greater than 0, such as 0.1, 0.2, and so on. Although the CP distribution is discrete, the independent variable of the distribution function is continuous, so it is expressed here using derivatives.
Constraint 1: (Guaranteed text diversity)
$DAI(x)\geq x,\forall x\in\left[\epsilon^{+},1\right).$ (5)
If the steganographic embedding can only select a token with extremely high CP, the generated diversity will be lost, and the existence of the stegos will be easily exposed. This constraint indicates that the number of codings allocated to tokens with lower CPs should be appropriately increased to enhance their selection likelihood.
Constraint 2: (Guaranteed text quality)
${\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}\geq\epsilon^{+},\forall x\in\left[\epsilon^{+},1\right),$ (6)
@@
The existing embedding ways assign the same number of codings to all tokens, which results in the same likelihood of randomly selecting each token, thus affecting the generation quality. Therefore, this constraint requires that tokens with larger CP need to be assigned a longer interval, that is, the number of allocations needs to increase with the increase of CP, and the change rate of the number of allocations cannot be too small. If ${\textstyle{{\partial\;DAI(x)}\over{\partial\;x}}}\leq\epsilon^{+}$, which means the number of allocation tokens is similar. It will degenerate into a similar allocation of the same number of codings as the existing work, thereby increasing the likelihood of selecting poor tokens.
Constraint 3: (Guaranteed to facilitate embedding)
${\frac{{\partial\ {DAI}{(x)}}}{{\partial\ x}}}\leq{\frac{{\partial\ x}}{{\partial\ x}}}\cap{\frac{{\partial^{2}{DAI}{(x)}}}{{\partial x^{2}}}}\leq 0,\forall x\in\left[\epsilon^{+},1\right).$ (7)
Steganography requires an algorithm that is conducive to embedding information. If you just want high-quality tokens to be selected and blindly increase the interval length of tokens with high CPs, the same prefix at the beginning and end of the interval will be shortened or even non-existent. This is not conducive to embedding secrets and deviates from the basic function of steganography. Therefore, in order to ensure the payload, it is necessary to avoid high CP tokens occupying too much interval length, and the growth trend of ${DAI}{(x)}$ should be gradually slow, that is, mathematically, it can be achieved by limiting the second-order derivative to be less than 0. When the conditions are equal, the corresponding function will be transformed into a linear function.
Here, we give several nonlinear functions:
@@
The smaller $\beta$ is, the smaller the interval of tokens with larger CP will be, while still satisfying the premise that the interval of tokens with larger CP is larger. It will increase the length of the prefix, while reducing the length of stegos to be generated, improving the payload.
Please note that these functions are not fixed, here are just a few examples that meet the constraints. These constraints are not exhaustive, and the final condensed expression is also for simplicity and easy to understand. This can just provide thinking for subsequent research that is more conducive to hiding information.
Loop the stego generation, and stego will be obtained. Algorithm 1 illustrates the embedding process.
@@
Secret extraction process
Bitstream extraction and embedding are a pair of inverse operations. Same as Discop [17], we also need to use the same scheme, initial context (instruction), and stegos. The instructions remain requirements visible to everyone in social networks. The formula of the extraction process $Ex$ is as follows:
$\left\{\begin{array}[]{l}Ex:\text{locate}\;\ \{{i_{n+1}}\;\ {\rm{in}}\;\ DAI(BC(P))\},\\ \quad\quad\;\text{match}\;\{i(R_{begin})\;\text{and}\;i(R_{end})\}\\ B_{j}=Ex(CPG(I)),B_{j}\in B\end{array}\right..$ (10)
The interval is located by the stego’s next token at this moment, and the same prefix is the currently extracted secret.